How any milliliters of 15% acid should be mixed with 150 milliliters of 40% acid to make 35 % mixture solution
A: write an equation
B: solve for x

Answer by stanbon(75887)

How any milliliters of 15% acid should be mixed with 150 milliliters of 40% acid to make 35 % mixture solution
--------------
Equation:
acid + acid = acid
0.40*150 + 0.15x = 0.35(150+x)
-------
40*150 + 15x = 35*150 + 35x
------
5*150 = 30x
x = 5*5 = 25 ml (amt. of 15% solution needed)
==================
Cheers,
Stan H.
==================






A: write an equation
B: solve for x